OpenStudy (anonymous):

will medal and fan just a few 8thgrade questions Solve the following system of equations. 8x + 5y = -105 3x + y = -14 x = -29, y = 5 x = 12, y = -50 x = 5, y = -40 x = 5, y = -29

OpenStudy (anonymous):

can you help me

OpenStudy (ghostgate):

I'm going to assume that your question is trying to solve for x and y with these two equations correct? 8x + 5y = -105 3x + y = -14

OpenStudy (anonymous):

yes

OpenStudy (ghostgate):

Okay, do you know the elimination method, I know there's an alternative way to solve for this, but the elimination method is the method I use the most.

OpenStudy (anonymous):

i'm not even supposed to be doing this yet so no my mom want me to be ahead e.e

OpenStudy (ghostgate):

Okay, with the elimination method, we take both equations and we have to even out one or the other variable, whichever seems smaller and easier to get rid of, for instance the 'y' variable. So... we'll want to multiply one of the equations by a number to even out the 'y' variables, so... what number should we use to multiply against the second equation to even out the 'y' variables for elimination? 8x + 5y = -105 3x + y = -14

OpenStudy (anonymous):

we should muliply 7?

OpenStudy (ghostgate):

No, okay let me give you a quick example to explain: 2x + 3y = 5 x + 4y = 10 Okay, so we would target the 'x' variable for elimination considering how small it is, so...: 2x + 3y = 5 2(x + 4y = 10) - Do you see what I did? --- 2x + 3y = 5 2x + 8y = 20 - Now that the variables are negative we subtract both equations. 2x - 2x = 0, 3y - 8y = -5y, 5 - 20 = -15 - Now we are left with: 0 - 5y = -15 - Since there is no 'x' variable we can now solve for 'y'. -5y = -15 - Divide. y = 3 - Do you get that? To find 'x', plug 'y' back into one of the equations and simply solve for it like any ordinary problem from before. Do you get that?

OpenStudy (anonymous):

yea you have to plug a piece of y equation into the x equation then solve it

OpenStudy (anonymous):

the timer went off

OpenStudy (ghostgate):

Alright, now for your problem I'm going to be a bit more technical: 8x + 5y = -105 3x + y = -14 ~ We target the 'y' variable and multiply the 2nd equation by '5'. --- 8x + 5y = -105 5(3x + y = -14) ~ Now we multiply 5 by the 2nd equation. --- 8x + 5y = -105 15x + 5y = -70 ~ Now that the 'y' variables are even, subtract the top equation by the bottom equation. --- 8x - 15x = -7x, 5y - 5y = 0, -105 + 70 = -35 ~ Now we have: -7x = -35 ~ Divide both sides by -7 to get your answer for x. -7x/-7 = -35/-7 = 1x = 5 ~ Which can be written as: x = 5 ~ Now to solve for 'y': --- Take one of the original equations, we'll take the 2nd one: 3x + y = -14 ~ Input your '5' for your 'x' variable. 3(5) + y = -14 ~ Multiply 3 by 5. 15 + y = -14 ~ Subtract both sides by 15. 15 - 15 + y = -14 - 15 ~ Which gives you: 0 + y = -29 ~ So you have: y = -29 ~ Check the equation: 3x + y = -14 ~ Input known variables. 3(5) + -29 = -14 ~ Multiply. 15 - 29 = -14 ~ Subtract. -14 = -14 ~ So the answer is correct. x = 5 y = -29 Do you get that?

OpenStudy (anonymous):

yea i'll try a couple of questions on my own using this later but the questionn timer went off i'm doing study island

OpenStudy (anonymous):

questions

OpenStudy (ghostgate):

I've never heard of Study Island, but... does that mean you can't do the questions or...?

OpenStudy (anonymous):

no we can still do questions but i can't do that one till it shows up again

OpenStudy (anonymous):

Which of the following relations is a function? (2, 4), (-4, 6), (2, 3), (-8, 2) (2, 0), (-4, 3), (8, 1), (-4, 5) (2, 4), (-4, 2), (8, 1), (-8, 2) (8, 1), (-4, 4), (2, 1), (8, 2) this is the next one

OpenStudy (anonymous):

gotta go for now i'll fan and medal so i can bother you later XD

OpenStudy (ghostgate):

Lol - Alright. ;)

OpenStudy (ghostgate):

The problem with your next question is that when it comes to slopes I'm not the best, and if I recall correctly since you have coordinates for a graph, that means it has something to do with a slope. Sorry that I can't help you there.

OpenStudy (anonymous):

i can help

OpenStudy (anonymous):

a function is when none of the X values repeat.